A small grinding wheel has a moment of inertia of 4.0*10^-5 kg*m2 . What net torque must be applied to the wheel for its angular acceleration to be 150rad/s2 ?
Answer #1

torque = moment of inertia*angular acceleration = (4*10-5)*150 = 6*10-3